from the edge of the top of the building with the height of 30 m erica kicks a soccer ball with a velocity of 12 m/s horizontally what is the velocity of the soccer ball when it strikes the ground
1
Expert's answer
2021-10-15T10:22:40-0400

The vertical component of velocity:


"v_y=\sqrt{2gh}=24.2\text{ m/s}."

The horizontal component does not change if we ignore air resistance. So, the velocity before hitting the ground can be found by the Pythagorean theorem:


"v=\sqrt{v_x^2+v_y^2}=27\text{ m/s}."